# Through the Dual-Prism: A Spectral Perspective on Graph Data Augmentation for Graph Classification

### v1

@@ 6 Conclusion
In this study, we adopt a spectral perspective, bridging graph properties and spectral insights for property-retentive and globally-aware graph data augmentation. Stemming from this point, we propose a novel augmentation method called Dual-Prism (DP), including DP-Noise and DP-Mask. By focusing on different frequency components in the spectrum, our method skillfully preserves graph properties while ensuring diversity in augmented graphs. Our extensive evaluations validate the efficacy of our methods across various learning paradigms on the graph classification task. In summary, our contributions highlight the potential of leveraging spectral insights in graph data augmentation.
Appendix A Details of Datasets
@@
We conduct our experiments on 21 different graph real-world datasets for graph classification tasks. In this section, we provide detailed descriptions of the datasets used in this paper. Specifically, for the supervised learning setting, we include a total of eight datasets from the TUDatasets benchmark (Morris et al., 2020) (i.e., PROTEINS, NCI1, IMDB-BINARY, IMDB-MULTI, REDDIT-BINARY, REDDIT-MULTI-5K, and REDDIT-MULTI-12K) and the OGB benchmark (Hu et al., 2020a) (i.e., ogbg-molhiv). For the semi-supervised learning setting, we include seven different datasets from the TUDatasets benchmark (Morris et al., 2020) (i.e., PROTEINS, NCI1, DD, COLLAB, GITHUB, REDDIT-BINARY, and REDDIT-MULTI-5K). For the unsupervised learning setting, we include seven different datasets from the TUDatasets benchmark (Morris et al., 2020) (i.e., PROTEINS, NCI1, DD, MUTAG, IMDB-BINARY, REDDIT-BINARY, and REDDIT-MULTI-5K). The detailed of these datasets can be found in Table 5.
@@
For the transfer learning setting, we pre-train on ZINC-2M chemical molecule dataset (Sterling & Irwin, 2015; Gómez-Bombarelli et al., 2018), and fine-tune on eignt different datasets, namely BBBP, Tox21, ToxCast, SIDER, ClinTox, MUV, HIV, and BACE. The detailed of these datasets can be found in Table 6.
@@ Appendix B Details of Baselines
  • DropEdge (Rong et al., 2019) selectively drops a portion of edges from the input graphs.
  • DropNode (Feng et al., 2020) omits a specific ratio of nodes from the provided graphs.
  • Subgraph (You et al., 2020) procures subgraphs from the main graphs using a random walk sampling technique.
  • M-Mixup (Verma et al., 2019) blends graph-level representations through linear interpolation.
  • SubMix (Yoo et al., 2022) combines random subgraphs from paired input graphs.
  • G-Mixup (Han et al., 2022) employs a class-focused graph mixup strategy by amalgamating graphons across various classes.
  • S-Mixup (Ling et al., 2023) adopts a mixup approach for graph classification, emphasizing soft alignments.
@@
  • GAE (Kipf & Welling, 2016b) is a non-probabilistic graph auto-encoder model, which is a variant of the VGAE (variational graph autoencoder).
  • Informax (Veličković et al., 2018) trains a node encoder to optimize the mutual information between individual node representations and a comprehensive global graph representation.
  • GraphCL (You et al., 2020) conducts an in-depth exploration of graph structure augmentations, including random edge removal, node dropping, and subgraph sampling.
@@
  • GL (graphlet kernel) (Pržulj, 2007) measures the similarity between graphs by counting the occurrences of small subgraphs, known as graphlets, within them. It captures local topological patterns, thus providing a comprehensive view of the graph structure.
  • WL (Weisfeiler-Lehman sub-tree kernel) (Shervashidze et al., 2011) captures the similarity between graphs by comparing subtrees of increasing heights. It effectively distinguishes non-isomorphic graphs and is often employed for graph classification tasks.
  • DGK (deep graph kernel) (Yanardag & Vishwanathan, 2015) combines the strengths of both graph kernels and deep learning. It leverages convolutional neural networks to learn hierarchical representations of graphs, enabling the kernel to capture complex patterns and structures within the data for a more refined similarity measure.
  • node2vec (Grover & Leskovec, 2016) captures low-dimensional embeddings of graph nodes by leveraging random walks originating from target nodes.
  • sub2vec (Adhikari et al., 2018) seeks to capture feature representations of arbitrary subgraphs, addressing the limitations of node-centric embeddings.
  • graph2vec (Narayanan et al., 2017) is a neural embedding framework designed to learn data-driven distributed representations of entire graphs.
  • InfoGraph (Sun et al., 2019) is designed to maximize the mutual information between complete graph representations and various substructures, such as nodes, edges, and triangles.
@@
  • MVGRL (Hassani & Khasahmadi, 2020) establishes a link between the local Laplacian matrix and a broader diffusion matrix by leveraging mutual information. This approach yields representations at both the node and graph levels, catering to distinct prediction tasks.
  • AD-GCL (Suresh et al., 2021) emphasize preventing the capture of redundant information during training. They achieve this by optimizing adversarial graph augmentation strategies in GCL and introducing a trainable non-i.i.d. edge-dropping graph augmentation.
  • JOAO (You et al., 2021) utilize a bi-level optimization framework to sift through optimal strategies, exploring multiple augmentation types like uniform edge or node dropping and subgraph sampling.
  • GCL-SPAN (Lin et al., 2022) introduces a spectral augmentation approach, which directs topology augmentations to maximize spectral shifts.
@@
  • EdgePred (Hamilton et al., 2017) employs an inductive approach that utilizes node features, such as text attributes, to produce node embeddings by aggregating features from a node’s local neighborhood, rather than training distinct embeddings for each node.
  • AttrMasking (Attribute Masking) (Hu et al., 2020b) is a pre-training method for GNNs that harnesses domain knowledge by discerning patterns in node or edge attributes across graph structures.
  • ContextPred (Context Prediction) (Hu et al., 2020b) is designed for pre-training GNNs that simultaneously learn local node-level and global graph-level representations via subgraphs to predict their surrounding graph structures.
@@
Appendix C Details of Experiments Settings
  • Experiment of Figure 3. In Figures 3a and 3b, labels in REDDIT-MULTI-12K for Class A and Class B are 1 and 10, respectively. The added edges in 3c are 3$\leftrightarrow$5, 3$\leftrightarrow$7, 1$\leftrightarrow$6, 2$\leftrightarrow$6, 0$\leftrightarrow$2, 1$\leftrightarrow$3, 1$\leftrightarrow$4, 2$\leftrightarrow$4, 4$\leftrightarrow$6, and 5$\leftrightarrow$7, respectively. The dropped edges in 3d are 0$\leftrightarrow$4, 3$\leftrightarrow$4, 4$\leftrightarrow$5, 4$\leftrightarrow$7, 0$\leftrightarrow$1, 2$\leftrightarrow$3, 0$\leftrightarrow$3, 5$\leftrightarrow$6, 6$\leftrightarrow$7, and 1$\leftrightarrow$2, respectively. The change of spectrum $\Delta L_{2}$ is the $L_{2}$ distance between the spectrum of $\mathcal{G}$ and augmented graph $\mathcal{G}^{\prime}$, denoted as $\Delta L_{2}=\sqrt{\sum_{i}(\lambda_{i}(\mathcal{G})-\lambda_{i}(\hat{\mathcal{G}}))^{2}}$, where $\lambda(\mathcal{G})$ represent the spectrum of $\mathcal{G}$ and $\lambda(\hat{\mathcal{G}})$ is the spectrum of $\hat{\mathcal{G}}$.
  • Experiment of Figure 5. For both Figure 5a and 5b, we employ GIN as the backbone model and conduct experiments over five runs. In Figure 5b, while testing one parameter, we draw the other parameters from their respective search spaces: $\sigma\in[0.1,0.5,1.0,2.0]$, $r_{f}\in[0.1,0.2,0.3,0.4,0.5,0.6,0.7,0.8]$, and $r_{a}\in[0,0.2,0.4,0.6,0.8,1]$. In addition, to control the noise adding to low- and high-frequency eigenvalues are in the same scale, the augmented $i$-th eigenvalue is calculated as $\lambda_{i}=\max(0,1+\epsilon)\times\lambda_{i}$, where $\epsilon\sim\mathcal{N}(0,\sigma)$.
@@ Appendix C Details of Experiments Settings
Supervised learning. Following prior works (Han et al., 2022; Ling et al., 2023), we utilize two GNN models, namely GCN (Kipf & Welling, 2017) and GIN (Xu et al., 2018). Details of these GNNs are provided below.
@@
For all other hyper-parameter search space and training configurations of the experiments on the IMDB-B, IMDB-M, REDD-B, REDD-M5, and REDD-M12, we keep consistent with Han et al. (2022). For all other hyper-parameter search space and training configurations of the experiments on the PROTEIN, NCI1, and ogbg-hiv, we keep consistent with Ling et al. (2023). Note that instead of adopting the results of baseline methods on the ogbg-hiv dataset from the reference directly, we reported the results of rerunning the baseline experiments on the ogbg-hiv dataset, which is higher than the results in the reference.
@@
Semi-supervised learning. We maintain consistency with You et al. (2020) for all hyper-parameter search spaces and training configurations. For all datasets, we conduct experiments at label rates of 1% (provided there are more than 10 samples for each class) and 10%. These experiments are performed five times, with each instance corresponding to a 10-fold evaluation. We report both the mean and standard deviation of the accuracies in percentages. During pre-training, we perform a grid search, tuning the learning rate among 0.01, 0.001, 0.0001 and the epoch number within 20, 40, 60, 80, 100.
@@
Unsupervised representation learning. Following You et al. (2020); Lin et al. (2022), we use a 5-layer GIN as encoders. For all hyper-parameter search spaces and training settings in unsupervised learning, we also align with the configurations presented in You et al. (2020). We conduct experiments five times, with each iteration corresponding to a 10-fold evaluation. The reported results in Table 3 include both the mean and standard deviation of the accuracy percentages.
@@
Transfer learning. Following the transfer learning setting in Hu et al. (2020b); You et al. (2020); Lin et al. (2022), we conduct graph classification experiments on a set of biological and chemical datasets via GIN models. Specifically, an encoder was first pre-trained on the large ZINC-2M chemical molecule dataset (Sterling & Irwin, 2015; Gómez-Bombarelli et al., 2018) and then was evaluated on small datasets from the same domains (i.e., BBBP, Tox21, ToxCast, SIDER, ClinTox, MUV, HIV, and BACE).
Appendix D More Experiments Results
@@
Empirical studies. In Section 3, we investigate spectral alterations due to adding an edge in a toy graph (depicted in Figure 2a). The spectral changes are presented in Figure 2b. Further, in Figure 7, we analyze the consequences on the spectrum when edges from the same toy graph are removed. Notably, the removal of edges 0-4 and 3-4 results in minimal spectral variations. However, the removal of edges 5-6 and 6-7 leads to pronounced changes in both high and low frequencies, evident from the pronounced shifts in values $\lambda_{2}$ and $\lambda_{5}$. To further illustrate our observation, we present an additional example featuring a nine-node toy graph that also exhibits similar results, as shown in Figure 8.
@@
Hyperparameter sensitivities. We conducted experiments on the IMDB-BINARY dataset, leveraging various combinations of standard deviation $\sigma$ and frequency ratio $r_{f}$ for both low and high-frequency components to assess the impacts of DP-Noise parameters. For these experiments, we employed the GIN as our backbone model let $\sigma$ and $r_{f}$ from two search spaces, where $\sigma\in[0.1,0.5,1.0,2.0]$ and $r_{f}\in[0.1,0.2,0.3,0.4,0.5,0.6,0.7,0.8]$. We run 5 experiments and report the average values in Figure 9. Our observations indicate that introducing noise in the high-frequency components tends to enhance the test accuracy more markedly than when infused in the low-frequency regions, as illustrated by the prevailing lighter color in Figure 9a. This observation resonates with the insights gleaned from Section 3. Building upon these general observations, we further elucidate the effects of each specific parameter below.
@@
Complexity and Time Analysis. Theoretically, the computational bottleneck of our data augmentation method primarily stems from the eigen-decomposition and reconstruction of the Laplacian matrix. For a graph with $n$ nodes, the computational complexity of both operations is $O(n^{3})$. In terms of implementation, we have measured the time cost required by our method. For each $n$, we randomly generated $100$ graphs and recorded the average time and standard deviation required for our data augmentation method. The results, presented in Table 8, are measured in milliseconds. The average number of nodes in commonly used graph classification datasets is approximately between $10$ and $500$. Therefore, in the majority of practical training scenarios, the average time consumption of our algorithm for augmenting a single graph is roughly between $1$ millisecond and $40$ milliseconds. The experiments conducted here did not employ any parallel computing or acceleration methods. However, in actual training processes, it is common to parallelize data preprocessing using multiple workers or to precompute and store the eigen-decomposition results of training data. Therefore, the actual time consumption required for our method in implementations will be even lower.
@@
Appendix E More Discussions
Intuitions & Advantages of proposed methods. (1) Properties Preservation. Data augmentation should not only increase the quantity of training data but also enrich the quality of the learning experience for the model. Here, quality refers to the diversity, relevance, and realism of the augmented data. Therefore, property-retentive augmentations provide a more genuine learning context for the model, thus directly improving performance. (2) Global Perspective. Looking at the graph globally allows us to understand the larger structures and patterns within the graph. By making broader changes to the graph’s structure, global augmentations can create more diverse training instances, compared to local augmentations which might only create minor variations of the existing instances, therefore enhancing understanding of complex graph relationships and contributing to improved performance.
@@ Appendix E More Discussions
Broader Impact. Through a spectral lens, our Dual-Prism (DP) augmentation method presents both significant advancements and implications in the realm of graph-based learning. This can lead to improved performance, robustness, and generalizability of graph neural networks (GNNs) across a myriad of applications, from social network analysis to molecular biology. In addition, by utilizing spectral properties, our method provides a more transparent approach to augmentation. This can help researchers and practitioners better understand how alterations to graph structures impact learning outcomes, thereby aiding in the interpretability of graph data augmentation.
@@
Limitations & Future Directions. A potential limitation of this study is its primary emphasis on homophily graphs. In contrast, heterophily graphs, where high-frequency information plays a more crucial role, are not extensively addressed (Bo et al., 2021). Looking ahead, it would be worth investigating learning strategies tailored to selectively alter eigenvalues, ensuring adaptability across diverse datasets. This includes developing methods to safely create realistic augmented graphs and experimenting with mix-up techniques involving eigenvalues from different graphs.
Comparison with Existing Works. There are two related works about the spectral view on graph data augmentation, i.e., (Liu et al., 2022; Lin et al., 2022), while both are grounded in the GCL framework. Specifically, Liu et al. (2022) proposes a rule to find the optimal contrastive pair under the GCL framework instead of a general augmentation method. GCL-SPAN Lin et al. (2022) centers on maximizing variance in the spectral domain, our observations indicate that overall spectral changes don’t always align with graph properties, as detailed in Section 3. Thus, constraining specific eigenvalues to remain invariant might be a more effective strategy for generating valid augmented graphs. Despite GCL-SPAN (Lin et al., 2022) also utilizing a spectral perspective, it in fact still modifies the spatial domain while optimizing in the spectral realm. In contrast, our techniques directly make alterations in the spectral domain, leading to more meaningful and effective alterations. This is evident by our methods’ superior performance in graph classification (see in Section 5.3) and underscores the efficiency of straightforward spectral modifications in creating more effective and discerning augmented graphs for graph classification. In addition, the DP method is specifically designed for graph classification tasks. Unlike node classification tasks Yoo et al. (2022) that emphasize node features and local structures, our approach is rooted in the analysis of global graph structures.

### v3

@@ Conclusion
In this study, we adopt a spectral perspective, bridging graph properties and spectral insights for property-retentive and globally-aware graph data augmentation. Stemming from this point, we propose a novel augmentation method DP, including DP-Noise and DP-Mask. By focusing on different frequency components in the spectrum, our method skillfully preserves graph properties while ensuring diversity in augmented graphs. Our extensive experiments validate the efficacy of our methods across various learning paradigms on the graph classification task.
Acknowledgements
This work is partially supported by the Advanced Research and Technology Innovation Centre (ARTIC), the National University of Singapore under Grant (project number: ELDT-RP2), and by National Research Foundation, Singapore, under its AI Singapore Programme (AISG Award No: AISG2-RP-2021-023). It is also supported by the National Natural Science Foundation of China (No. 62402414). Xavier Bresson is supported by NUS Grant ID R-252-000-B97-133.
Appendix
Details of Datasets
@@ Acknowledgements
We conduct our experiments on 21 different graph real-world datasets for graph classification tasks. In this section, we provide detailed descriptions of the datasets used in this paper. Specifically, for the supervised learning setting, we include a total of eight datasets from the TUDatasets benchmark (Morris et al. 2020) (i.e., PROTEINS, NCI1, IMDB-BINARY, IMDB-MULTI, REDDIT-BINARY, REDDIT-MULTI-5K, and REDDIT-MULTI-12K) and the OGB benchmark (Hu et al. 2020a) (i.e., ogbg-molhiv). For the semi-supervised learning setting, we include seven different datasets from the TUDatasets benchmark (Morris et al. 2020) (i.e., PROTEINS, NCI1, DD, COLLAB, GITHUB, REDDIT-BINARY, and REDDIT-MULTI-5K). For the unsupervised learning setting, we include seven different datasets from the TUDatasets benchmark (Morris et al. 2020) (i.e., PROTEINS, NCI1, DD, MUTAG, IMDB-BINARY, REDDIT-BINARY, and REDDIT-MULTI-5K). The detailed of these datasets can be found in Table 5.
@@
For the transfer learning setting, we pre-train on ZINC-2M chemical molecule dataset (Sterling and Irwin 2015; Gómez-Bombarelli et al. 2018), and fine-tune on eignt different datasets, namely BBBP, Tox21, ToxCast, SIDER, ClinTox, MUV, HIV, and BACE. The detailed of these datasets can be found in Table 6.
@@
  • DropEdge (Rong et al. 2019) selectively drops a portion of edges from the input graphs.
  • DropNode (Feng et al. 2020) omits a specific ratio of nodes from the provided graphs.
  • Subgraph (You et al. 2020) procures subgraphs from the main graphs using a random walk sampling technique.
  • M-Mixup (Verma et al. 2019) blends graph-level representations through linear interpolation.
  • SubMix (Yoo, Shim, and Kang 2022) combines random subgraphs from paired input graphs.
  • G-Mixup (Han et al. 2022) employs a class-focused graph mixup strategy by amalgamating graphons across various classes.
  • S-Mixup (Ling et al. 2023) adopts a mixup approach for graph classification, emphasizing soft alignments.
@@
  • GAE (Kipf and Welling 2016b) is a non-probabilistic graph auto-encoder model, which is a variant of the VGAE (variational graph autoencoder).
  • Informax (Veličković et al. 2018) trains a node encoder to optimize the mutual information between individual node representations and a comprehensive global graph representation.
  • GraphCL (You et al. 2020) conducts an in-depth exploration of graph structure augmentations, including random edge removal, node dropping, and subgraph sampling.
@@
  • sub2vec (Adhikari et al. 2018) seeks to capture feature representations of arbitrary subgraphs, addressing the limitations of node-centric embeddings.
  • graph2vec (Narayanan et al. 2017) is a neural embedding framework designed to learn data-driven distributed representations of entire graphs.
  • InfoGraph (Sun et al. 2019) is designed to maximize the mutual information between complete graph representations and various substructures, such as nodes, edges, and triangles.
@@
  • MVGRL (Hassani and Khasahmadi 2020) establishes a link between the local Laplacian matrix and a broader diffusion matrix by leveraging mutual information. This approach yields representations at both the node and graph levels, catering to distinct prediction tasks.
  • AD-GCL (Suresh et al. 2021) emphasize preventing the capture of redundant information during training. They achieve this by optimizing adversarial graph augmentation strategies in GCL and introducing a trainable non-i.i.d. edge-dropping graph augmentation.
  • JOAO (You et al. 2021) utilize a bi-level optimization framework to sift through optimal strategies, exploring multiple augmentation types like uniform edge or node dropping and subgraph sampling.
  • GCL-SPAN (Lin, Chen, and Wang 2022) introduces a spectral augmentation approach, which directs topology augmentations to maximize spectral shifts.
@@
  • EdgePred (Hamilton, Ying, and Leskovec 2017) employs an inductive approach that utilizes node features, such as text attributes, to produce node embeddings by aggregating features from a node’s local neighborhood, rather than training distinct embeddings for each node.
  • AttrMasking (Attribute Masking) (Hu et al. 2020b) is a pre-training method for GNNs that harnesses domain knowledge by discerning patterns in node or edge attributes across graph structures.
  • ContextPred (Context Prediction) (Hu et al. 2020b) is designed for pre-training GNNs that simultaneously learn local node-level and global graph-level representations via subgraphs to predict their surrounding graph structures.
@@
Details of Experiments Settings
  • Experiment of Figure 4. In Figures 4a and 4b, labels in REDDIT-MULTI-12K for Class A and Class B are 1 and 10, respectively. The added edges in 4c are 3$\leftrightarrow$5, 3$\leftrightarrow$7, 1$\leftrightarrow$6, 2$\leftrightarrow$6, 0$\leftrightarrow$2, 1$\leftrightarrow$3, 1$\leftrightarrow$4, 2$\leftrightarrow$4, 4$\leftrightarrow$6, and 5$\leftrightarrow$7, respectively. The dropped edges in 4d are 0$\leftrightarrow$4, 3$\leftrightarrow$4, 4$\leftrightarrow$5, 4$\leftrightarrow$7, 0$\leftrightarrow$1, 2$\leftrightarrow$3, 0$\leftrightarrow$3, 5$\leftrightarrow$6, 6$\leftrightarrow$7, and 1$\leftrightarrow$2, respectively. The change of spectrum $\Delta L_{2}$ is the $L_{2}$ distance between the spectrum of $\mathcal{G}$ and augmented graph $\mathcal{G}^{\prime}$, denoted as $\Delta L_{2}=\sqrt{\sum_{i}(\lambda_{i}(\mathcal{G})-\lambda_{i}(\hat{\mathcal{G}}))^{2}}$, where $\lambda(\mathcal{G})$ represent the spectrum of $\mathcal{G}$ and $\lambda(\hat{\mathcal{G}})$ is the spectrum of $\hat{\mathcal{G}}$.
  • Experiment of Figure 3. For both Figure 3a and 3b, we employ GIN as the backbone model and conduct experiments over five runs. In Figure 3b, while testing one parameter, we draw the other parameters from their respective search spaces: $\sigma\in[0.1,0.5,1.0,2.0]$, $r_{f}\in[0.1,0.2,0.3,0.4,0.5,0.6,0.7,0.8]$, and $r_{a}\in[0,0.2,0.4,0.6,0.8,1]$. In addition, to control the noise adding to low- and high-frequency eigenvalues are in the same scale, the augmented $i$-th eigenvalue is calculated as $\lambda_{i}=\max(0,1+\epsilon)\times\lambda_{i}$, where $\epsilon\sim\mathcal{N}(0,\sigma)$.
Supervised learning. Following prior works (Han et al. 2022; Ling et al. 2023), we utilize two GNN models, namely GCN (Kipf and Welling 2017) and GIN (Xu et al. 2018). Details of these GNNs are provided below.
@@
For all other hyper-parameter search space and training configurations of the experiments on the IMDB-B, IMDB-M, REDD-B, REDD-M5, and REDD-M12, we keep consistent with (Han et al. 2022). For all other hyper-parameter search space and training configurations of the experiments on the PROTEIN, NCI1, and ogbg-hiv, we keep consistent with (Ling et al. 2023). Note that instead of adopting the results of baseline methods on the ogbg-hiv dataset from the reference directly, we reported the results of rerunning the baseline experiments on the ogbg-hiv dataset, which is higher than the results in the reference.
@@
Semi-supervised learning. We maintain consistency with (You et al. 2020) for all hyper-parameter search spaces and training configurations. For all datasets, we conduct experiments at label rates of 1% (provided there are more than 10 samples for each class) and 10%. These experiments are performed five times, with each instance corresponding to a 10-fold evaluation. We report both the mean and standard deviation of the accuracies in percentages. During pre-training, we perform a grid search, tuning the learning rate among 0.01, 0.001, 0.0001 and the epoch number within 20, 40, 60, 80, 100.
@@
Unsupervised representation learning. Following (You et al. 2020; Lin, Chen, and Wang 2022), we use a 5-layer GIN as encoders. For all hyper-parameter search spaces and training settings in unsupervised learning, we also align with the configurations presented in (You et al. 2020). We conduct experiments five times, with each iteration corresponding to a 10-fold evaluation. The reported results in Table 4 include both the mean and standard deviation of the accuracy percentages.
@@
Transfer learning. Following the transfer learning setting in (Hu et al. 2020b; You et al. 2020; Lin, Chen, and Wang 2022), we conduct graph classification experiments on a set of biological and chemical datasets via GIN models. Specifically, an encoder was first pre-trained on the large ZINC-2M chemical molecule dataset (Sterling and Irwin 2015; Gómez-Bombarelli et al. 2018) and then was evaluated on small datasets from the same domains (i.e., BBBP, Tox21, ToxCast, SIDER, ClinTox, MUV, HIV, and BACE).
More Experiments Results
Transfer learning. We draw comparisons between our methods and six baselines, including a reference model without pre-training (referred to No-Pre-Train), Informax (Veličković et al. 2018), EdgePred (Hamilton, Ying, and Leskovec 2017), AttrMasking (Hu et al. 2020b), ContextPred (Hu et al. 2020b) and GraphCL (You et al. 2020). The comparative results are shown in Table 8. Our methods demonstrated SOTA performance, outperforming competitors on half of the datasets. Especially, our methods consistently outperform the conventional GraphCL method, which indicates our data augmentation methods as better choices for graph contrastive learning. Notably, DP-Mask achieves an 83.52% ROC-AUC score on ClinTox, exceeding the performance of GraphCL by a substantial margin (nearly 10%). These findings demonstrate the enhanced efficacy of our techniques in the transfer learning setting for graph classification tasks.
Table 8: Performance comparisons in the transfer learning setting. The best and second best results are highlighted with bold and underline, respectively. The metric is ROC-AUC scores (%). * and ** denote the improvement over the second-best baseline is statistically significant at levels 0.1 and 0.05, respectively. Baseline results are taken from (Hu et al. 2020b; You et al. 2020).
Dataset BBBP Tox21 ToxCast SIDER ClinTox MUV HIV BACE
No-Pre-Train 65.80$\pm$ 4.50 74.00$\pm$ 0.80 63.40$\pm$ 0.60 57.30$\pm$ 1.60 58.00$\pm$ 4.40 71.80$\pm$ 2.50 75.30$\pm$ 1.90 70.10$\pm$ 5.40
Infomax 68.80$\pm$ 0.80 75.30$\pm$ 0.50 62.70$\pm$ 0.40 58.40$\pm$ 0.80 69.90$\pm$ 3.00 75.30$\pm$ 2.50 76.00$\pm$ 0.70 75.90$\pm$ 1.60
EdgePred 67.30$\pm$ 2.40 76.00$\pm$ 0.60 64.10$\pm$ 0.60 60.40$\pm$ 0.70 64.10$\pm$ 3.70 74.10$\pm$ 2.10 76.30$\pm$ 1.00 79.90$\pm$ 0.90
AttrMasking 64.30$\pm$ 2.80 76.70$\pm$ 0.40 64.20$\pm$ 0.50 61.00$\pm$ 0.70 71.80$\pm$ 4.10 74.70$\pm$ 1.40 77.20$\pm$ 1.10 79.30$\pm$ 1.60
ContextPred 68.00$\pm$ 2.00 75.70$\pm$ 0.70 63.90$\pm$ 0.60 60.90$\pm$ 0.60 65.90$\pm$ 3.80 75.80$\pm$ 1.70 77.30$\pm$ 1.00 79.60$\pm$ 1.20
GraphCL 69.68$\pm$ 0.67 73.87$\pm$ 0.66 62.40$\pm$ 0.57 60.53$\pm$ 0.88 75.99$\pm$ 2.65 69.80$\pm$ 2.66 78.47$\pm$ 1.22 75.38$\pm$ 1.44
DP-Noise 70.38$\pm$ 0.91* 74.33$\pm$ 0.42 64.08$\pm$ 0.25 61.52$\pm$ 0.79 76.26$\pm$ 1.68 74.09$\pm$ 2.30 78.63$\pm$ 0.37 77.37$\pm$ 1.76
DP-Mask 71.63$\pm$ 1.86** 75.51$\pm$ 0.26 63.77$\pm$ 0.22 61.33$\pm$ 0.17 83.52$\pm$ 1.07** 75.02$\pm$ 0.36 78.73$\pm$ 0.77 79.45$\pm$ 0.69
@@
Empirical studies. In Section A Spectral Lens on Graph Data Augmentations, we investigate spectral alterations due to adding an edge in a toy graph (depicted in Figure 2a). The spectral changes are presented in Figure 2b. Further, in Figure 7, we analyze the consequences on the spectrum when edges from the same toy graph are removed. Notably, the removal of edges 0-4 and 3-4 results in minimal spectral variations. However, the removal of edges 5-6 and 6-7 leads to pronounced changes in both high and low frequencies, evident from the pronounced shifts in values $\lambda_{2}$ and $\lambda_{5}$. To further illustrate our observation, we present an additional example featuring a nine-node toy graph that also exhibits similar results, as shown in Figure 8.
@@
Hyperparameter sensitivities. We conducted experiments on the IMDB-BINARY dataset, leveraging various combinations of standard deviation $\sigma$ and frequency ratio $r_{f}$ for both low and high-frequency components to assess the impacts of DP-Noise parameters. For these experiments, we employed the GIN as our backbone model let $\sigma$ and $r_{f}$ from two search spaces, where $\sigma\in[0.1,0.5,1.0,2.0]$ and $r_{f}\in[0.1,0.2,0.3,0.4,0.5,0.6,0.7,0.8]$. We run 5 experiments and report the average values in Figure 9. Our observations indicate that introducing noise in the high-frequency components tends to enhance the test accuracy more markedly than when infused in the low-frequency regions, as illustrated by the prevailing lighter color in Figure 9a. This observation resonates with the insights gleaned from Section A Spectral Lens on Graph Data Augmentations. Building upon these general observations, we further elucidate the effects of each specific parameter below.
@@
Complexity and Time Analysis. Theoretically, the computational bottleneck of our data augmentation method primarily stems from the eigen-decomposition and reconstruction of the Laplacian matrix. For a graph with $n$ nodes, the computational complexity of both operations is $O(n^{3})$. In terms of implementation, we have measured the time cost required by our method. For each $n$, we randomly generated $100$ graphs and recorded the average time and standard deviation required for our data augmentation method. The results, presented in Table 9, are measured in milliseconds. The average number of nodes in commonly used graph classification datasets is approximately between $10$ and $500$. Therefore, in the majority of practical training scenarios, the average time consumption of our algorithm for augmenting a single graph is roughly between $1$ millisecond and $40$ milliseconds. The experiments conducted here did not employ any parallel computing or acceleration methods. However, in actual training processes, it is common to parallelize data preprocessing using multiple workers or to precompute and store the eigen-decomposition results of training data. Therefore, the actual time consumption required for our method in implementations will be even lower. Therefore, although the complexity is theoretically high, in fact, the time required is practically low.
@@
More Discussions
Advantages of the proposed method. At the heart of data augmentation lies a dual objective: to preserve the essence of the data while expanding its diversity. Our approach strikes this balance skillfully. Leveraging spectral view, it safeguards key graph property (Figure 1e), ensuring the augmented data remains authentic and relevant. This not only enhances the training data’s quality but also directly boosts the performance, as detailed in the following section.
Comparison with existing works. Our method distinguishes itself from three related spectral view works: (Liu et al. 2022), (Lin, Chen, and Wang 2022), and (Yang et al. 2023). (Liu et al. 2022) focuses on optimal contrastive pair selection instead of a general augmentation method, and GCL-SPAN (Lin, Chen, and Wang 2022) centers on maximizing spectral variance. However, our observations suggest that broad spectral changes do not always correspond with graph properties. In addition, though GCL-SPAN also use a spectral perspective, it essentially alters the spatial domain, instead of direct spectral modifications, while our method’s enhanced performance underlines the effectiveness of spectral alterations. GASSER(Yang et al. 2023) perturbs different frequency bands to uphold homophily ratios and fine-tune task-specific features. Whereas our approach concentrates on the essential structural integrity required for graph classification.
@@
Broader Impact. Through a spectral lens, our proposed augmentation method presents both significant advancements and implications in the realm of graph-based learning. This can lead to improved performance, robustness, and generalizability of graph neural networks (GNNs) across a myriad of applications, from social network analysis to molecular biology. In addition, by utilizing spectral properties, our method provides a more transparent approach to augmentation. This can help researchers and practitioners better understand how alterations to graph structures impact learning outcomes, thereby aiding in the interpretability of graph data augmentation. While there is a superficial overlap in terms of using ”spectrum” for ”graph data augmentation” with some related works, e.g., (Liu et al. 2022; Lin, Chen, and Wang 2022), the core methodologies and outcomes are distinctly different. We believe it should be a new direction and open up new avenues for spectral-based augmentation techniques, allowing more diverse research on doing the augmentation on the spectrum.
@@
Limitations & Future Directions. A potential limitation of this study is its primary emphasis on homophily graphs. In contrast, heterophily graphs, where high-frequency information plays a more crucial role, are not extensively addressed (Bo et al. 2021). In addition, as we have discussed on the characteristics of molecular datasets above, in some situations, the high-frequency component may also include important information for graph classification tasks. Looking ahead, it would be worth investigating learning strategies tailored to selectively alter eigenvalues, ensuring adaptability across diverse datasets. This includes developing methods to safely create realistic augmented graphs and experimenting with mix-up techniques involving eigenvalues from different graphs.
High-Frequency Significance in Molecular Data. Incorporating insights from Table 8, our method shows varied performance across datasets. Specifically, within the Tox21 and ToxCast toxicology datasets, our approach does not achieve top-ranking results. This could be due to the potential significance of high-frequency components in capturing essential structural details within these datasets, which aligns with the intricate biological nature of the targets and stress response pathways represented in these datasets. The Tox21 dataset contains qualitative toxicity measurements for various biological targets, such as nuclear receptors and stress response pathways(Ma et al. 2022), which could be notably influenced by subtle molecular changes detected in the high-frequency domain. Consequently, it is plausible to suggest that high-frequency components play a vital role in distinguishing between toxic and non-toxic molecules by capturing subtle molecular differences crucial for predicting bioactivity and toxicity. In contrast to datasets focusing on broader properties like BBBP (Blood-Brain Barrier permeability) or BACE (inhibitor activity against beta-secretase), Tox21 and ToxCast encompass a wide range of bioassays that may be particularly sensitive to the high-frequency intricacies of molecular structures. Thus, the complexity of molecular compositions and the diversity of bioactivity assays in Tox21 and ToxCast imply that high-frequency components may contain essential information for toxicity prediction that may not be as pronounced in datasets such as BBBP.